At the beginning of the year, Kevin Company’s liabilities equaled P60,000. During the year, assets increased by P80,000, and at year-end assets equaled P180,000. Liabilities decreased by P10,000 during the year. What are the beginning and ending amounts of equity?



1
Expert's answer
2022-01-17T17:04:03-0500

Assets= equity+liabilities

Beginning of the year

L=P60000

A=P100000

E=?

A=E+L

100000=E+60000

Beginning of the year

Equity=40000

End of the year equity?

L=(60000-100000)=50000

Closing assets =P180000

180000=E+50000

180000-50000=E

End year equity=P130000